OpenStudy (anonymous):

solve the equation show all work must be in simplest form. x+9/11=8/11

OpenStudy (katrinakaif):

x = -1 why Cross Mulitply giving: 88 = 11(x + 9) Distribut ethe 11 88 = 11x + 99 Substract 99 from both sides 88-99 = 11x -11 = 11x Divide by 11 -1 = x

OpenStudy (gw2011):

If you mean the problem is x + (9/11) = 8/11, then it becomes: x + (9/11) = 8/11 x = 8/11 - 9/11 x = -1/11